# Rethinking the Evaluation of Dialogue Systems: Effects of User Feedback on Crowdworkers and LLMs

### v1

@@ 6. Significance of User Feedback
In this section, we examine the impact of user feedback from follow-up utterances on reducing annotator variability in their evaluation labels as part of our analysis to answer RQ2. To assess agreement, we calculate the standard deviation of workers’ scores for each turn and categorize the data into two groups: Group 1 (scores below the median standard deviation, indicating high agreement) and Group 2 (scores above the median, indicating low agreement). We find that interestingness and explanation quality consistently exhibit higher agreement among annotators when the system response is uninteresting or lacks explanation. However, there is no clear agreement pattern among workers for relevance and usefulness.
@@
This analysis identifies specific scenarios where user feedback plays a pivotal role, addressing ambiguous requests by providing clarity, making generic requests more specific and actionable, simplifying complex requests, and compensating for annotators’ lack of domain knowledge. In these scenarios, user feedback consistently improves the overall quality and consistency of the annotation process, highlighting its significance in enhancing system evaluations.
For instance, consider a user who requested a movie recommendation, noting their fondness for ‘A Nightmare on Elm Street (1984)’. This request could imply a preference for several aspects: the classic horror genre, the directorial style of Wes Craven, or the specific era of the 1980s. Initially, the system recommends a popular horror-comedy from 2017, which, while relevant in terms of genre, might not align precisely with the nuanced preferences implied in the user’s request. The initial ratings from the crowdworkers reflect this broad interpretation, with Worker A giving a score of 70/100 based on the assumption of a preference for any modern horror film, Worker B scoring 55/100 considering the genre match but not the era, and Worker C giving 60/100, focusing on the horror aspect but overlooking the director and the time period. However, the user’s follow-up utterance, stating a specific enjoyment for Wes Craven’s horror style and a preference for 1980s movies, brings new clarity. With this additional information, the crowdworkers in Setup 2 assess the same recommendation against these more specific criteria. Worker D now scores 40/100, recognizing the mismatch in both director and era, Worker E assigns 30/100, acknowledging the recommendation’s lack of alignment with the specific director and the time period, and Worker F scores 35/100, understanding that while the genre is somewhat aligned, the specific preferences for Wes Craven’s style and the 1980s era are not met.
This significant change in ratings, from relatively high to low in Setup 2, illustrates that without user feedback, crowdworkers may overestimate the usefulness of a system’s response based on a broader, less nuanced interpretation of the request. However, when provided with more detailed preferences from the user, their ratings become more aligned and tend to be lower, reflecting a more accurate assessment of how well the system’s response meets the user’s refined criteria. This highlights the importance of the user’s next utterance in accurately gauging the usefulness of system responses, especially in the context of complex requests.
@@ 7. Sources and Bias
In this section we answer RQ3. To understand the basis of workers’ assessments and their choices when assigning evaluation labels, we conducted a survey where we asked workers to indicate the sources of information they relied on when making judgments. Both setups offered the same options for information sources, except for follow-up utterance, which was available only in Setup 2. The available sources included personal knowledge, searched online, guessed, user request, and system response. Additionally, we performed a manual analysis of the workers’ justifications to evaluate any potential biases introduced by their chosen information sources.
@@
Figure 6. Distribution of sources workers relied on to make their judgment for the four aspects in the two setups.
@@
A small number of workers utilize online sources, primarily when assessing relevance and usefulness. This implies that workers without domain knowledge leverage online information for more accurate assessments. Interestingly, there is a decrease in the percentage of workers using online sources in Setup 2, specifically for evaluating usefulness. This suggests that the introduction of the user’s follow-up utterance in Setup 2 acts as an additional information source, assisting workers lacking domain knowledge in assessing the system response’s usefulness.
@@ 8. Discussion and Conclusion
Relevance. We considered two experimental setups, one without the user’s follow-up utterance and one with. In both setups, crowdworkers and LLM-based annotators largely concur when evaluating relevance. However, subtle differences emerge with the inclusion of follow-up utterances, particularly for LLMs which tend to assign lower scores compared to humans. Though there is no significant mean difference in Setup 1 for both annotator groups, we note that LLM scores show a higher correlation to expert user satisfaction ratings than humans. Crowdworkers relied more on the system response, user request, and their prior knowledge to gauge the relevance of the recommended item. With lots of candidate movies to recommend, crowdworkers may lack knowledge of some of these movies to assess their relevance, which results into making an educated guess. However, compared to humans LLMs are rich in internal knowledge on these movies, thus improving their relevance assessment.
@@
Interestingness. The aspect of interestingness presented unique challenges. Both crowdworkers and LLMs exhibited lower correlations with user satisfaction ratings, indicating that both annotator groups struggled to capture the user’s subjective perception of interestingness. The presence of user feedback had a limited impact on improving assessments in this aspect. This is also observed with the low Kappa and ICC scores in Table 1. Nonetheless, utterances such as “that’s interesting,” “sounds good,” and “you are funny” lead to an increase in annotator agreement and correlation with user satisfaction ratings (Setup 2), emphasizing the significance of user feedback in improving system evaluation.
@@
In general, there is a distinct difference in ratings assigned by both annotator groups in the two setups, indicating that user feedback does influence system evaluation. Human workers are more susceptible to user feedback in usefulness and interestingness compared to LLMs in interestingness. User feedback leads to personalized usefulness assessment by workers and improves worker agreement when uncertainty arises in the user’s request. The lack of adaptability to user feedback by the LLM in assessing usefulness, suggests that LLMs may require additional mechanisms such as prompt tuning to enhance user-centric evaluations by LLMs. However, user feedback can sometimes lead to assessments that do not align with overall user satisfaction, resulting in lower correlation scores. It is important to note that this study employed a single LLM for annotation, and results may vary with different LLMs. Additionally, the study did not consider potential biases in the crowdworker pool or LLM training data, which could impact the results. Finally, the findings may not be generalizable to all conversational systems, and further research is needed to validate these insights in diverse contexts.
Acknowledgements.
This research was supported by the Dreams Lab, a collaboration between Huawei Finland, the University of Amsterdam, and the Vrije Universiteit Amsterdam, by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, by project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), and by the FINDHR (Fairness and Intersectional Non-Discrimination in Human Recommendation) project that received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.

### v2

@@ 6. Significance of User Feedback
In this section, we examine the impact of user feedback from follow-up utterances on reducing annotator variability in their evaluation labels as part of our analysis to answer (RQ2). To assess agreement, we calculate the standard deviation of workers’ scores for each turn and categorize the data into two groups: Group 1 (scores below the median standard deviation, indicating high agreement) and Group 2 (scores above the median, indicating low agreement). We find that interestingness and explanation quality consistently exhibit higher agreement among annotators when the system response is uninteresting or lacks explanation. However, there is no clear agreement pattern among workers for relevance and usefulness.
@@
This analysis identifies specific scenarios where user feedback plays a pivotal role, such as addressing ambiguous requests by providing clarity, making generic requests more specific and actionable, simplifying complex requests, and compensating for annotators’ lack of domain knowledge. In these scenarios, user feedback consistently improves the overall quality and consistency of the annotation process, highlighting its significance in enhancing system evaluations. [^4] A detailed example of a complex user request can be found in our Online Appendix in Sec. B.1.
@@ 7. Sources and Bias
In this section, we answer (RQ3). To understand the basis of workers’ assessments and their choices when assigning evaluation labels, we conducted a survey where we asked workers to indicate the sources of information they relied on when making judgments. Both setups offered the same options for information sources, except for follow-up utterance, which was available only in Setup 2. The available sources included personal knowledge, searched online, guessed, user request, and system response. Additionally, we performed a manual analysis of the workers’ justifications to evaluate any potential biases introduced by their chosen information sources.
@@
Figure 6. Distribution of sources workers relied on to make their judgments for the four aspects in the two setups.
@@
A few workers utilize online sources, primarily when assessing relevance and usefulness, implying that workers without domain knowledge leverage online information for more accurate assessments. Interestingly, there is a decrease in the percentage of workers using online sources in Setup 2, specifically for evaluating usefulness. This suggests that the introduction of the user’s follow-up utterance in Setup 2 acts as an additional information source, assisting workers lacking domain knowledge in assessing the system response’s usefulness.
@@ 8. Discussion and Conclusion
Relevance. We considered two experimental setups, one without the user’s follow-up utterance and one with. In both setups, crowdworkers and LLM-based annotators largely concur when evaluating relevance. However, subtle differences emerge with the inclusion of follow-up utterances, particularly for LLMs which tend to assign lower scores compared to humans. Though there is no significant mean difference in Setup 1 for both annotator groups, we note that LLM scores show a higher correlation to expert user satisfaction ratings than humans. Crowdworkers relied more on the system response, user request, and their prior knowledge to gauge the relevance of the recommended item. With lots of candidate movies to recommend, crowdworkers may lack knowledge of some of these movies to assess their relevance, which results in making an educated guess. However, compared to humans LLMs are rich in internal knowledge on these movies, thus improving their relevance assessment.
@@
Interestingness. The aspect of interestingness presented unique challenges. Both crowdworkers and LLMs exhibited lower correlations with user satisfaction ratings, indicating that both annotator groups struggled to capture the user’s subjective perception of interestingness. The presence of user feedback had a limited impact on improving assessments in this aspect. This is also observed with the low Kappa and ICC scores in Tab. 1. Nonetheless, utterances such as “that’s interesting,” “sounds good,” and “you are funny” lead to an increase in annotator agreement and correlation with user satisfaction ratings (Setup 2), emphasizing the significance of user feedback in improving system evaluation.
@@
In general, there is a distinct difference in ratings assigned by both annotator groups in the two setups, indicating that user feedback does influence system evaluation. Human workers are more susceptible to user feedback in usefulness and interestingness compared to LLMs in interestingness. User feedback leads to personalized usefulness assessment by workers and improves worker agreement when uncertainty arises in the user’s request. The lack of adaptability to user feedback by the LLM in assessing usefulness, suggests that LLMs may require additional mechanisms such as prompt tuning to enhance user-centric evaluations by LLMs. Therefore, it is important to assign annotation tasks to LLMs based on the nature of the task, leveraging their strengths in objective assessments like relevance annotation while complementing with human assessors for tasks demanding subjective evaluation or sensitivity to user preferences and feedback. Combining human annotators and LLMs can lead to better system evaluations by leveraging the unique strengths of each type of annotator for specific evaluation tasks.
However, user feedback can sometimes lead to assessments that do not align with overall user satisfaction, resulting in lower correlation scores as observed for the relevance aspect in Setup 2. It is important to note that this study employed a single LLM for annotation, and results may vary with different LLMs. Additionally, potential biases in the crowdworker pool and the LLM’s training data could influence the findings. For future work, we will conduct further research to validate our findings across diverse conversational systems.
Online Appendix
Further analyses and discussions are available at https://github.com/Clemenciah/LLMCrowdDialogueEval/tree/main/Appendix.
Data
We publicly release the annotated data at https://github.com/Clemenciah/LLMCrowdDialogueEval/tree/main/Data.
Acknowledgments
This research was supported by the Dreams Lab, a collaboration between Huawei Finland, the University of Amsterdam, and the Vrije Universiteit Amsterdam, by the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and by the European Union’s Horizon Europe program under grant agreement No 101070212.
All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.